10. Suppose (X, Y) is uniformly distributed over the disk 2 + y 36. Then
Answer #1

here P(0<X<2 and 0 <y<3) =(area covered with square of area from 0<X<2 and 0 <y<3) /area of total disk =(2*3)/(\pi*62) =6/(36*\pi) =1/(6\pi)